OpenStudy (anonymous):

I need help finding the limit of t-t(sqrt(t))/2t^3/2 + 3t - 5 as x approaches infinity. I think I'm doing something wrong algebraically. The answer I came up with is 1/2. I think this answer is wrong. Please help by breaking down the steps so I can check my work. Thank you. :)

OpenStudy (anonymous):

Are you just considering the powers of the numerator and denominator?

OpenStudy (anonymous):

\[\lim_{t\to\infty}\frac{t-t\sqrt t}{2t^{3/2}+3t-5}=\lim_{t\to\infty}\frac{t\color{red}{-t^{3/2}}}{\color{red}{2t^{3/2}}+3t-5}\]